Circular No. 7811

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ap IN M74 K. Kinugasa and H. Kawakita, Gunma Astronomical Observatory

(GAO); K. Ayani and T. Kawabata, Bisei Astronomical Observatory (BAO); and H. Yamaoka, Kyushu University, write: “Low-resolution spectra of SN 2002ap (IAUC 7810) were obtained on Jan. 31.4 UT with the GAO 0.65-m telescope (+ GCS; range 380–750 nm) and on Jan. 31.5 with the BAO 1.01-m telescope (range 470–700 nm). Preliminary reduction reveals a rather blue continuum with a steep decrease over 650–700 nm, without any deep absorption or emission. Very broad (FWZI ∼ 30–50 nm) and shallow depressions exist around 470, 570 (deepest), and 620 nm. The overall features resemble that of the peculiar type-Ib/c supernova (or ‘hypernova’) 1997ef, but SN 2002ap is much bluer.”

P. Meikle and L. Lucy, Imperial College, London; S. Smartt, Institute of Astronomy, Cambridge; B. Leibundgut, European Southern Observatory; and P. Lundqvist, Stockholm Observatory, report that R. Ostensen obtained a spectrum (range 432.6–789.9 nm) of SN 2002ap at the William Herschel Telescope on Jan. 30.9 UT: “The spectrum shows a blue continuum with broad absorption features at about 448.0, 557.5, and 719.5 nm. The red absorption is particularly broad, spanning ≈ 660–760 nm. The absorption features bear some resemblance to those seen in SN 1998bw at 7 days (Stathakis et al. 2000, MNRAS 314, 807), though shifted further to the blue in SN 2002ap by 15 000–20 000 km/s. The 448.0-nm absorption appears to be deeper and may include a more narrow component than in SN 1998bw. The emission peak at 530 nm for SN 1998bw is absent in the spectrum of SN 2002ap. We suggest that SN 2002ap may be a younger and/or more energetic version of SN 1998bw.”

A. Gal-Yam and O. Shemmer, Tel Aviv University, report: “A spectrum (range 400–780 nm) of SN 2002ap was obtained with the Wise Observatory 1-m telescope (+ FOSC spectrograph) on Jan. 31 by J. Dann. The object was very bright, and the high S/N spectrum shows a very blue continuum with broad features around 400 and 500 nm, a steep decline (of about a factor of 3 in flux) from 500 nm to 700 nm, and a broad hump around 600 nm superposed on the blue continuum. A hint of a rise toward the red end of the spectrum, at 780 nm, is also evident. Overall, the spectrum is markedly similar to spectra of the peculiar type-Ic supernova 1998bw obtained before maximum light (e.g., Patat et al. 2001, Ap.J. 555, 900). In view of the peculiar properties of this object, as well as the possible connection of SN 1998bw with GRB 980425, we urge further observations at all wavelengths.”

2002 January 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green